# Event-Grounding Graph: Unified Spatio-Temporal Scene Graph from Robotic Observations

### v2

@@ Abstract
A fundamental aspect for building intelligent autonomous robots that can assist humans in their daily lives is the construction of rich environmental representations. While advances in semantic scene representations have enriched robotic scene understanding, current approaches lack a connection between spatial features and dynamic events; e.g., connecting the blue mug to the event washing a mug. In this work, we introduce the  event-grounding graph (EGG), a framework grounding event interactions to spatial features of a scene. This representation allows robots to perceive, reason, and respond to complex spatio-temporal queries. Experiments using real robotic data demonstrate EGG’s capability to retrieve relevant information and respond accurately to human inquiries concerning the environment and events within. Furthermore, the EGG framework’s source code and evaluation dataset are released as open-source at: https://github.com/aalto-intelligent-robotics/EGG.
@@ V-B Task-based graph pruning
Finally, we text-serialize $\mathcal{G}_{Q}$ into JSON format, which is then integrated into the prompt for the LLM agent to respond to the user’s query [^1] All prompts and specific LLM used can be found in the paper codebase..
@@ VI-A Experimental setup
The robot task involves verbalizing its observations of a scene through different forms of scene representation. We manually label each video with ground truth summary captions and object role captions to assess the capability of VLMs in comparison to human judgment. For each event, we also manually provide segmentation masks for each object at their first and last appearances, enabling projection to capture their instance positions and incorporate them into the position history of their respective spatial nodes (refer to Sec. V-A for implementation details of the spatial graph). The dataset comprises 80 questions with ground truth answers across four modalities, for instance [^2] The full dataset, including the events and question-answer pairs, will be made publicly available after acceptance to preserve anonymity.:
@@ IX Acknowledgement
The authors acknowledge the use of the MIDAS infrastructure of Aalto School of Electrical Engineering. The authors also acknowledge the use of Grammarly’s generative AI tool and OpenAI’s GPT4o [19] for improving readability across all sections. Additionally, the authors acknowledge the role of OpenAI’s GPT4o [19] in evaluating our method as detailed in Sec. VI, as well as generating documentation and improving code readability in the project’s code base.

### v3

@@ Abstract
A fundamental aspect for building intelligent autonomous robots that can assist humans in their daily lives is the construction of rich environmental representations. While advances in semantic scene representations have enriched robotic scene understanding, current approaches lack a connection between spatial features and dynamic events; e.g., connecting the blue mug to the event washing a mug. In this work, we introduce  event-grounding graph (EGG), a framework grounding event interactions to spatial features of a scene. This representation allows robots to perceive, reason, and respond to complex spatio-temporal queries. Experiments using real robotic data demonstrate EGG’s capability to retrieve relevant information and respond accurately to human inquiries concerning the environment and events within. Furthermore, the EGG framework’s source code and evaluation dataset are released as open-source at: https://github.com/aalto-intelligent-robotics/EGG.
@@ V-B Task-based graph pruning
Finally, we text-serialize $\mathcal{G}_{Q}$ into JSON format, which is then integrated into the prompt for the LLM agent to respond to the user’s query [^1] All prompts and specific LLM used can be found in the paper’s codebase..
@@ VI-A Experimental setup
The robot task involves verbalizing its observations of a scene through different forms of scene representation. We manually label each video with ground truth summary captions and object role captions to assess the capability of VLMs in comparison to human judgment. For each event, we also manually provide segmentation masks for each object at their first and last appearances, enabling projection to capture their instance positions and incorporate them into the position history of their respective spatial nodes (refer to Sec. V-A for implementation details of the spatial graph). The dataset comprises 80 questions with ground truth answers across four modalities, for instance [^2] The full dataset, including the events and question-answer pairs, will be made publicly available.:
@@ IX Acknowledgement
The authors acknowledge the use of Grammarly’s generative AI tool and OpenAI’s GPT4o [19] for improving readability across all sections. Additionally, the authors acknowledge the role of OpenAI’s GPT4o [19] in evaluating our method as detailed in Sec. VI, as well as generating documentation and improving code readability in the project’s code base.